evaluate the following integral
∫ sec^2 (x) / tan x + 2 dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{sec^{2}x}{tanx+2}dx
tanx + 2 = t
sec^{2}xdx = dt
I = \int \frac{dt}{t}
I =log(t) + c
I = log(tanx + 2) + c